# UniMapGen: A Generative Framework for Large-Scale Map Construction from Multi-modal Data

### v1

@@ 4.2 Ablation Study
To validate the effectiveness of key components, as shown in Tab. 2 and 3, we conduct ablation studies on OpenSatMap20 val set to explore the effectiveness of state update, line reorder, augmentation, the assistance of multi-modal inputs.
@@ 5 Limitation and Discussion
Although UniMapGen demonstrates promising results in large-scale map construction, it has several limitations. Firstly, the computational demands of MLLMs, along with their relatively slow inference speed (average 4 seconds per $\times135m$), pose practical challenges for real-time map construction like MapTR for autonomous driving. Actually, UniMapGen serves as a large-scale map constructor to assist autonomous driving and navigation, and could utilize up-to-date PV frames to update existing maps. Secondly, the text prompt in UniMapGen only supports coordinates for target map construction, but not more flexible natural language, such as generating the map from the school to the hospital. This is due to the lack of training data. Using an offline object detection model (Carion et al. 2020) to extract the object location or synthesizing the training data leveraging more annotations has the potential to solve this, which will be explored in our future work.
@@ Appendix A Model Architecture
BEV Encoder. We employ ViT-Large-14 from Dinov2 (Liu et al. 2024) as our pretrained BEV encoder, encoding a single satellite image $I^{BEV}$ into N D-dimensional tokens $\boldsymbol{e}^{BEV}\in\mathbb{R}^{D\times N}$. Dinov2 is trained with a self-supervised learning strategy on a diverse set of 1.2B images, including ImageNet-22k (Deng et al. 2009), the train split of ImageNet-1k (Deng et al. 2009), Google Landmarks (Weyand et al. 2020), and several fine-grained datasets. Please refer to  (Oquab et al. 2023) for more details about Dinov2.
PV Encoder. Ground-level frames provide more detailed, complete, and up-to-date road surface information, complementing satellite image’s limitations such as occlusions and outdated data. We utilize a 3D convolution (Tran et al. 2015) followed by the image encoder from Qwen2-VL-2B (Wang et al. 2024) (Qwen2-VL-ViT) as our PV Encoder, processing L PV frames $I^{PV}$ into M D-dimensional tokens $\boldsymbol{e}^{PV}\in\mathbb{R}^{D\times M}$. Qwen2-VL-ViT is a self-implemented ViT by the Qwen team with 675 parameters, which could receive input images at any resolution.
Vector Tokenizer and Detokenizer. We propose a special vector tokenizer to convert the previous state map $G_{n-1}$ vector sets into special tokens $\boldsymbol{e}^{G}$, subsequently processed through word embeddings, as mentioned in Sec 3.2 Map Serialization and Model Architecture. Tab. 5 illustrates an example. Specifically, each coordinate number is encoded into special tokens as a whole rather than split into single digits. For example, the number 257 will be tokenized into <257> rather than <2><5><7>. Additionally, the word will be tokenized as a whole as well, such as converting ’points’ to <points>. Finally, we remove all the spaces in the token string to reduce the number of tokens. On the contrary, the vector detokenizer performs the inverse operation, decoding the MLLM-predicted tokens (like the second line) back into vectorized maps (like the first line).
Format
Vector {’points’: [[257,49],…,[376,15]],’category’:’Curb’,…}
Tokens
<{><points><:><[><[><257><,><49><]>,…,<[><376>
<,><15><]><]><,><category><:><Curb>,…<}>
Large Language Model and Text Encoder. We adopt Qwen2.5-1.5B (Qwen et al. 2025) as our LLM, receiving multimodal input to construct maps. Additionally, we utilize the LLM’s text tokenizer to tokenize $T^{prompt}$ into $\boldsymbol{e}^{T}$. Qwen2.5-1.5B is a dense, decoder-only language model, featuring 1.54 billion parameters and designed for efficient instruction-following tasks. Its architecture incorporates advanced components such as Rotary Position Embeddings (RoPE), SwiGLU activations, RMSNorm normalization, and attention mechanisms with QKV bias and tied word embeddings. The model comprises 28 layers with 12 attention heads for the query and 2 for the key and value, supporting a context length of up to 32,768 tokens and a generation length of 8,192 tokens. The training process of Qwen2.5-1.5B involved two main stages:
@@ Appendix B Training Dataset Creation
BEV Only Training. We leverage the level-20 parts of OpenSatMap (Zhao et al. 2024) dataset, denoted as OpenSatMap20, as our BEV images. We use level-20 instead of level-19 because of its higher resolution at 0.15m/pixel. OpenSatMap20 dataset includes 1180 training and 393 validation satellite image of 4096*4096 pixels. Each image in the dataset is annotated with lane-level vectorized maps, with each line classified into three line categories (Curb, Lane line, and Virtual Line), and eight attributes (e.g., line types and level of occlusion). Following (Zhao et al. 2024), we only use line categories (Curb, Lane line, and Virtual Line) and line types (solid line, thick solid line, dashed line, short dashed line and others) as our line attributes of training data. To meet the training data requirement for MLLMs, we conduct overlapped and inclined cropping data augmentation (as illustrated in Sec. C), creating nearly 700k 896*896 training patches.
PV Only Training. As illustrated in (Zhao et al. 2024), the satellite images in OpenSatMap cover the cities in nuSences and Argoverse2. We leverage the PV frames from nuSences (Caesar et al. 2020) and Argoverse2 (Wilson et al. 2023) datasets, enabling map construction from PV frames. To ensure the consistency of training data, we use the modified annotation in OpenSatMap dataset. To be specific, we first use the GPS information provided by nuSences, Argoverse2, and OpenSatMap dataset to link the PV frames to the corresponding BEV images and their annotations. Then, similar to MapTR, we crop the lines within 60m height and 30m width as ground truth, considering the perception field of PV frames. The PV images are resized into 644*364.
@@
  • Automatic processing. This aims to create the training data of text prompts for the targeted map paired with the corresponding targeted map vectors. To achieve this, we utilize the Segment Anything Model (SAM) (Kirillov et al. 2023) to generate precise road area segmentations for each line vector annotation in OpenSatMap. Specifically, we input the annotated points from line vectors directly into SAM, which effectively segments the corresponding road regions. To identify complete road segments and their associated line vectors, we employ a region merging algorithm. Specifically, we merge segmented regions with an Intersection over Union (IoU) threshold of 0.5, while simultaneously combining their corresponding line vector annotations. This process results in comprehensive road segments paired with their complete set of line vector representations.
@@
Appendix C Data Augmentation
As the training images of OpenSatMap are quite limited (1180 at 4096*4096). If they are cropped to 896*896 resolution without augmentation, there are only 18880 patches, which is hard to meet the large-scale training data request of Multi-modal Large Language Model. The risk of overfitting the MLLM is heightened. To mitigate this, as shown in Fig. 4, we apply an overlapped and inclined cropping data augmentation to increase the number of training patches from 2K to 700K. The i,j-th cropped patch is formatted as:
$patch_{i,j}=\text{Crop}(I^{BEV},C_{i,j},\theta),$ (6)
Appendix D Lane Topology Construction
@@
Appendix E Parameters
Training Parameters. We trained our UniMapGen for 6 epochs with a batch size of 32 and the fully fine-tuning strategy applied. We used BF16 mixed precision training with DeepSpeed (Rasley et al. 2020) to increase the training efficiency. We used the AdamW (Loshchilov and Hutter 2017) optimizer and a weight decay of 0.1. A cosine learning rate decay with a peak learning rate of 2e-5 and a linear warm-up of 100 steps was adopted.
@@ Appendix F Visualizations
Complex Examples. Fig. 6 shows several complex examples generated by UniMapGen from 4 different cities.
@@
Bad Cases. Fig. 7 illustrates several bad cases generated by UniMapGen. Although UniMapGen has inference ability in map construction, when the lines are extremely worn out or occluded without context (like where the line disappears and reappears among the tall trees), UniMapGen will be confused and construct inaccurate maps.

### v2

@@ 4.2 Ablation Study
To validate the effectiveness of key components, as shown in Tab. 3 and 4, we conduct ablation studies on OpenSatMap20 val set to explore the effectiveness of state update, line reorder, augmentation, the assistance of multi-modal inputs.
@@ 5 Limitation and Discussion
Although UniMapGen demonstrates promising results in large-scale map construction, it has several limitations. Firstly, the computational demands of MLLMs, along with their relatively slow inference speed (average 4 seconds per $\times135m$), pose practical challenges for real-time map construction like MapTR for autonomous driving. Actually, UniMapGen serves as a large-scale map constructor to assist autonomous driving and navigation, and could utilize up-to-date PV frames to update existing maps. Secondly, the text prompt in UniMapGen only supports coordinates for target map construction, but not more flexible natural language, such as generating the map from the school to the hospital. This is due to the lack of training data. Using an offline object detection model (carion2020end) to extract the object location or synthesizing the training data leveraging more annotations has the potential to solve this, which will be explored in our future work.
Although UniMapGen has inference ability in map construction, as shown in 6, when the lines are extremely worn out or occluded without context (like where the line disappears and reappears among the tall trees), UniMapGen will be confused and construct inaccurate maps.
@@ Appendix A Model Architecture
BEV Encoder. We employ ViT-Large-14 from Dinov2 (remoteclip) as our pretrained BEV encoder, encoding a single satellite image $I^{BEV}$ into N D-dimensional tokens $\boldsymbol{e}^{BEV}\in\mathbb{R}^{D\times N}$. Dinov2 is trained with a self-supervised learning strategy on a diverse set of 1.2B images, including ImageNet-22k (deng2009imagenet), the train split of ImageNet-1k (deng2009imagenet), Google Landmarks (weyand2020google), and several fine-grained datasets. Please refer to  (oquab2023dinov2) for more details about Dinov2.
PV Encoder. Ground-level frames provide more detailed, complete, and up-to-date road surface information, complementing satellite image’s limitations such as occlusions and outdated data. We utilize a 3D convolution (tran2015learning) followed by the image encoder from Qwen2-VL-2B (wang2024qwen2vlenhancingvisionlanguagemodels) (Qwen2-VL-ViT) as our PV Encoder, processing L PV frames $I^{PV}$ into M D-dimensional tokens $\boldsymbol{e}^{PV}\in\mathbb{R}^{D\times M}$. Qwen2-VL-ViT is a self-implemented ViT by the Qwen team with 675 parameters, which could receive input images at any resolution.
Vector Tokenizer and Detokenizer. We propose a special vector tokenizer to convert the previous state map $G_{n-1}$ vector sets into special tokens $\boldsymbol{e}^{G}$, subsequently processed through word embeddings, as mentioned in Sec 3.2 Map Serialization and Model Architecture. Specifically, each coordinate number is encoded into special tokens as a whole rather than split into single digits. For example, the number 257 will be tokenized into <257> rather than <2><5><7>. Additionally, the word will be tokenized as a whole as well, such as converting ’points’ to <points>. Finally, we remove all the spaces in the token string to reduce the number of tokens. On the contrary, the vector detokenizer performs the inverse operation, decoding the MLLM-predicted tokens (like the second line) back into vectorized maps (like the first line).
Large Language Model and Text Encoder. We adopt Qwen2.5-1.5B (qwen2025qwen25technicalreport) as our LLM, receiving multimodal input to construct maps. Additionally, we utilize the LLM’s text tokenizer to tokenize $T^{prompt}$ into $\boldsymbol{e}^{T}$. Qwen2.5-1.5B is a dense, decoder-only language model, featuring 1.54 billion parameters and designed for efficient instruction-following tasks. Its architecture incorporates advanced components such as Rotary Position Embeddings (RoPE), SwiGLU activations, RMSNorm normalization, and attention mechanisms with QKV bias and tied word embeddings. The model comprises 28 layers with 12 attention heads for the query and 2 for the key and value, supporting a context length of up to 32,768 tokens and a generation length of 8,192 tokens. The training process of Qwen2.5-1.5B involved two main stages:
@@ Appendix B Training Dataset Creation
BEV Only Training. We leverage the level-20 parts of OpenSatMap (zhao2024opensatmap) dataset, denoted as OpenSatMap20, as our BEV images. We use level-20 instead of level-19 because of its higher resolution at 0.15m/pixel. OpenSatMap20 dataset includes 1180 training and 393 validation satellite image of 4096*4096 pixels. Each image in the dataset is annotated with lane-level vectorized maps, with each line classified into three line categories (Curb, Lane line, and Virtual Line), and eight attributes (e.g., line types and level of occlusion). Following (zhao2024opensatmap), we only use line categories (Curb, Lane line, and Virtual Line) and line types (solid line, thick solid line, dashed line, short dashed line and others) as our line attributes of training data. To meet the training data requirement for MLLMs, we conduct overlapped and inclined cropping data augmentation, creating nearly 700k 896*896 training patches.
PV Only Training. As illustrated in (zhao2024opensatmap), the satellite images in OpenSatMap cover the cities in nuSences and Argoverse2. We leverage the PV frames from nuSences (caesar2020nuscenes) and Argoverse2 (wilson2023argoverse) datasets, enabling map construction from PV frames. To ensure the consistency of training data, we use the modified annotation in OpenSatMap dataset. To be specific, we first use the GPS information provided by nuSences, Argoverse2, and OpenSatMap dataset to link the PV frames to the corresponding BEV images and their annotations. Then, similar to MapTR, we crop the lines within 60m height and 30m width as ground truth, considering the perception field of PV frames. The PV images are resized into 644*364.
@@
  • Automatic processing. This aims to create the training data of text prompts for the targeted map paired with the corresponding targeted map vectors. To achieve this, we utilize the Segment Anything Model (SAM) (kirillov2023segment) to generate precise road area segmentations for each line vector annotation in OpenSatMap. Specifically, we input the annotated points from line vectors directly into SAM, which effectively segments the corresponding road regions. To identify complete road segments and their associated line vectors, we employ a region merging algorithm. Specifically, we merge segmented regions with an Intersection over Union (IoU) threshold of 0.5, while simultaneously combining their corresponding line vector annotations. This process results in comprehensive road segments paired with their complete set of line vector representations.
@@
Appendix C Lane Topology Construction
@@
Appendix D Parameters
Training Parameters. We trained our UniMapGen for 6 epochs with a batch size of 32 and the fully fine-tuning strategy applied. We used BF16 mixed precision training with DeepSpeed (rasley2020deepspeed) to increase the training efficiency. We used the AdamW (loshchilov2017decoupled) optimizer and a weight decay of 0.1. A cosine learning rate decay with a peak learning rate of 2e-5 and a linear warm-up of 100 steps was adopted.